Tim is 4 years older than one-third of Juan's age. In 10 years, Juan will be 6 years older than twice as old as Tim is now. How old is each now?
Answer by ptaylor(2198)

Let x= Juan's current age; x+10=Juan's age in 10 years
Then (1/3)x+4=Tim's current age;
Now we are told that:
x+10=2{(1/3)x+4}+6 get rid of brackets
x+10=(2/3)x+8+6 =(2/3)x+14 multiply each side by 3
3x+30=2x+42 subtract 30 and also 2x from each side
3x-2x+30-30=2x-2x+42-30 collect like terms
x=12----------------Juan's current age
(1/3)x+4=(1/3)*12+4=8 -------Tim's current age


CK
---4 years older than 1/3 of Juan's age=(1/3)*12+4=8
8=8
---In 10 years, Juan will be 6 years older than twice as old as Tim is now.
12+10=22
2*8+6=22
22=22
Hope this helps---ptaylor